# LiDPM: Rethinking Point Diffusion for Lidar Scene Completion

### v1

@@ III-B LiDPM
The network $\epsilon_{\theta}$ has the architecture proposed in LiDiff, but with all batch normalization layers replaced by instance normalization. Indeed, in this architecture, $\epsilon_{\theta}$ takes as input widely different point clouds for conditioning: $\mathbf{p}^{s}$ or $\mathbf{0}$. We noticed that the statistics of the features may vary significantly when passing $\mathbf{p}^{s}$ or $\mathbf{0}$, potentially corrupting the running statistics and leading to instabilities during inference. To avoid this issue, we use instance normalization layers in which the statistics of the features are computed for each instance during training and inference. We elaborate on this in Sec. IV-D.
@@ IV-A Comparison to existing methods
TABLE I: SemanticKITTI validation set. Baselines, metrics and ground-truth data are from LiDiff [9].
@@
LMSCNet [37] Voxel - 0.431 32.2 23.1 3.5 0.641
LODE [18] Surface - 0.451 43.6 47.9 6.1 1.029
MID [2] Surface - 0.470 45.0 41.0 17.0 0.503
PVD [11] Points - 0.498 21.2 8.0 1.4 1.256
LiDiff [9] Points 0.564 0.444 42.5 33.3 11.1 0.434
LiDPM (ours) Points 0.532 0.440 45.5 43.9 16.1 0.446
LiDiff [9] Points 0.573 0.416 40.7 38.9 24.8 0.376
LiDPM   (ours) Points 0.542 0.403 44.4 44.0 27.6 0.377
@@
TABLE II: Ablation studies.
@@
50 0.437 43.9 45.1 19.4
20 0.428 44.9 44.8 18.2
10 0.443 41.7 43.2 20.7
5 0.447 41.2 42.1 20.9
Diffusion only. We first evaluate the quality of our learned global diffusion model, LiDPM. We compare it to the local diffusion model, LiDiff, and to PVD [11], whose results on SemanticKITTI were obtained by LiDiff authors [9]. No refinement is used after diffusion for any of these methods. The results are reported in Tab. I. We notice that LiDPM performs better than the other diffusion-only methods on all metrics, except for the Chamfer distance (CD), where it is just $1$ cm away from LiDiff. We also remark that our method’s IoU for voxels of size 0.5 m and 0.2 m is already better than the IoU obtained by LiDiff after refinement.
Complete method. We then compare to all baselines, still in Tab. I. We include the refinement step used for densification in LiDiff and LiDPM. We notice that LiDPM outperforms all the others in terms of JSD BEV and IoU at 0.2 m and 0.1 m. LiDPM is also better than LiDiff in JSD 3D and IoU at 0.5 m, and reaches an equivalent CD. We only notice a slight performance drop for LiDPM before and after refinement in JSD 3D and IoU at 0.5 m. The qualitative results presented in Fig. 4 show that the generated point clouds with LiDPM exhibit less spurious structures than those generated with LiDiff, in particular far away from the center of the scenes.
@@ IV-B Hyperparameter study
Start of the diffusion $t_{0}$. We generate point clouds using Eq. 2 for different starting points $t_{0}$, as defined in Eq. 4, and not using the DPM-Solver. We show in Tab. II(a) that the best results are observed for $t_{0}$ around $300$. Starting the diffusion with $t_{0}=50$ prevents the network from completing the point cloud (low recall), as the starting point cloud $\tilde{\mathbf{p}}^{d}_{t_{0}}$ is not noisy enough to fill all the gaps present in $\mathbf{p}^{s}$. In contrast, picking $t_{0}$ close to $T$ improves the recall, but the original structures in $\mathbf{p}^{s}$ are less preserved (lower precision). Starting at $t_{0}=300$ is a sweet spot to optimize the Chamfer distance and the voxel IoU. The conclusion remains the same after post-processing the results with the refinement network.
@@
DPM-Solver. For $t_{0}=300$, we study the achievable speedup to generate a point cloud by using the DPM-Solver. We test the number of steps to generate the point clouds, between $5$ and $50$ (see Tab. II(b)). We notice that $20$ steps are enough while the results degrade when going below.
@@ Appendix A Appendix
We present here the formulation of the diffusion process as a local denoising, as discussed and used in prior work [9, 31]. We then highlight our concerns, arguing that this approach can be more complex and may introduce approximations in the denoising process.
@@ A-A Formulation of the diffusion process as a local denoising
where $\lambda>0$, and $\mathcal{L}_{\rm mean},\mathcal{L}_{\rm std}$ return the mean and standard deviation of the entries in $\epsilon_{\theta}(\mathbf{p}^{d}_{t},t)$. When setting $\lambda=0$, the LiDiff authors noticed that the mean and standard deviation of $\epsilon_{\theta}$ were far from 00 and $1$, while $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Thus, the regularization loss $\mathcal{L}_{\rm reg}$ was introduced to guide the outputs of $\epsilon_{\theta}$ towards these expected values.
@@ A-B Limitations of local denoising paradigm
While the above local point diffusion proved to work, it holds on major approximations, which we outline here.
a. The justification of LiDiff to propose local point diffusion is that DDPM requires normalization, which leads to compression of data along some of the point cloud axes (Sec. 3.3 of LiDiff [9]). We believe this observation may be flawed. Not only could normalization be applied within fixed bounds, but it is actually not needed for DDPM, which can be formulated in the metric space (see Sec. III-B).
b. A major limitation of local diffusion is the reliance on a noisy estimate $\tilde{\mathbf{p}}^{s}$ of the ground truth in Eq. 9 and Eq. 12, which hinders its use for data generation, thereby losing one of the major properties of diffusion models.
c. Third, diffusion requires the predicted noise to converge to a zero-centered distribution. To compensate for the predicted local noise not being zero-centered, the authors add an additional regularization term $\mathcal{L}_{\rm reg}$ (Sec. 3.4 of LiDiff [9]), which further complexifies the optimization objective.
In this work, we show that DDPM can be applied on non-normalized point clouds without losing details after generation. Hence, the local (re-)formulation is not required (observation a.). Moreover, since we follow the vanilla ‘scene-level’ diffusion, our formulation does not require any additional regularization (observation b.). Last, a byproduct of our choices is that our simpler formulation opens up the possibility to generate point clouds (observation c.).

### v2

@@ III-B LiDPM
The network $\epsilon_{\theta}$ has the architecture proposed in LiDiff, but with all batch normalization layers replaced by instance normalization. Indeed, in LiDiff $\epsilon_{\theta}$ takes as input widely different point clouds for conditioning: $\mathbf{p}^{s}$ or $\mathbf{0}$. We noticed that the features statistics may vary significantly when passing $\mathbf{p}^{s}$ or $\mathbf{0}$, potentially corrupting the running statistics and leading to instabilities during inference. To avoid this issue, we use instance normalization layers in which the statistics of the features are computed for each instance during training and inference. We elaborate on this in Sec. IV-D.
@@ IV-A Comparison to existing methods
TABLE I: SemanticKITTI validation set. Baselines, metrics and ground-truth data are from LiDiff [9] [^1] The Voxel IoU values differ from the camera-ready version as per the correction in LiDiff evaluation code..
@@
LMSCNet [37] Voxel - 0.431 30.8 12.1 3.7 0.641
LODE [18] Surface - 0.451 33.8 16.4 5.0 1.029
MID [2] Surface - 0.470 31.6 22.7 13.1 0.503
PVD [11] Points - 0.498 15.9 4.0 0.6 1.256
LiDiff [9] Points 0.564 0.444 31.5 16.8 4.7 0.434
LiDPM (ours) Points 0.532 0.440 34.1 19.5 6.3 0.446
LiDiff [9] Points 0.573 0.416 32,4 23.0 13.4 0.376
LiDPM   (ours) Points 0.542 0.403 36.6 25.8 14.9 0.377
@@
TABLE II: Ablation studies.11footnotemark: 1
@@
50 0.437 32.9 19.2 6.8
20 0.428 33.9 19.6 6.5
10 0.443 31.0 18.0 6.9
5 0.447 30.7 17.5 6.8
Diffusion only. We first evaluate the quality of our learned global diffusion model, LiDPM. We compare it to the local diffusion model, LiDiff, and to PVD [11], whose results on SemanticKITTI were obtained by LiDiff authors [9]. No refinement is used after diffusion for any of these methods. The results are reported in Footnote 1. We notice that LiDPM performs better than the other diffusion-only methods on all metrics, except for the Chamfer distance (CD), where it is just $1$ cm away from LiDiff. We also remark that our method’s IoU for voxels of size 0.5 m and 0.2 m is already better than the IoU obtained by LiDiff after refinement.
Complete method. We then compare to all baselines, still in Footnote 1. We include the refinement step used for densification in LiDiff and LiDPM. We notice that LiDPM outperforms all the others in terms of JSD BEV and IoU at 0.2 m and 0.1 m. LiDPM is also better than LiDiff in JSD 3D and IoU at 0.5 m, and reaches an equivalent CD. We only notice a slight performance drop for LiDPM before and after refinement in JSD 3D and IoU at 0.5 m. The qualitative results presented in Fig. 4 show that the generated point clouds with LiDPM exhibit less spurious structures than those generated with LiDiff, in particular far away from the center of the scenes.
@@ IV-B Hyperparameter study
Start of the diffusion $t_{0}$. We generate point clouds using Eq. 2 for different starting points $t_{0}$, as defined in Eq. 4, and not using the DPM-Solver. We show in Footnote 1(a) that the best results are observed for $t_{0}$ around $300$. Starting the diffusion with $t_{0}=50$ prevents the network from completing the point cloud (low recall), as the starting point cloud $\tilde{\mathbf{p}}^{d}_{t_{0}}$ is not noisy enough to fill all the gaps present in $\mathbf{p}^{s}$. In contrast, picking $t_{0}$ close to $T$ improves the recall, but the original structures in $\mathbf{p}^{s}$ are less preserved (lower precision). Starting at $t_{0}=300$ is a sweet spot to optimize the Chamfer distance and the voxel IoU. The conclusion remains the same after post-processing the results with the refinement network.
@@
DPM-Solver. For $t_{0}=300$, we study the achievable speedup to generate a point cloud by using the DPM-Solver. We test the number of steps to generate the point clouds, between $5$ and $50$ (see Footnote 1(b)). We notice that $20$ steps are enough while the results degrade when going below.
@@ Appendix A Appendix
We present here the diffusion formulation as a local problem, as discussed and used in prior work [9, 31]. We then highlight concerns that this formulation adds complexity and may introduce approximations in the denoising process.
@@ A-A Formulation of the diffusion process as a local denoising
where $\lambda>0$, and $\mathcal{L}_{\rm mean},\mathcal{L}_{\rm std}$ guide the mean and standard deviation over $\epsilon_{\theta}(\mathbf{p}^{d}_{t},t)$ to 00 and $1$, respectively, as $\epsilon_{\theta}$ predicts peaky distribution, far from expected $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@ A-B Limitations of local denoising paradigm
While the above local point diffusion proved to work, it relies on major approximations, which we outline here.
a. The justification of LiDiff to propose local point diffusion is that DDPM requires normalization, which leads to compression of data along some of the point cloud axes (Sec. 3.3 of LiDiff). We believe this observation may be flawed. Not only could normalization be applied within fixed bounds, but it is actually not needed for DDPM, which can be formulated in the metric space (see Sec. III-B).
b. Diffusion requires the predicted noise to follow a standard normal distribution. To compensate for the predicted local noise deviating from a zero-mean, unit-variance Gaussian, the authors add regularization losses (Sec. 3.4 of LiDiff), which further complicates the optimization objective.
c. A key limitation of local diffusion is the reliance on a noisy estimate $\tilde{\mathbf{p}}^{s}$ of the ground truth in Eq. 9 and Eq. 12, which hinders its use for data generation, thereby losing one of the central benefits of diffusion models.
In this work, we show that DDPM can be applied to non-normalized point clouds without losing detail; hence, the local (re-)formulation is not required (a.). Moreover, our vanilla scene-level formulation avoids the need for added regularization (b.). Last, a byproduct of our choices opens up the possibility to generate point clouds (c.).